I need to know step by step how to find the value of (g o h)(2)
g(x) =5x2-2x+3 and h(x) √9x+7 find the value of (g o h )(2)

Answer by stanbon(75887)

find the value of (g o h)(2)
g(x) = 5x^2-2x+3 and h(x)= √9x+7
------------
goh(2) = g[h(2)] = g[sqrt(9*2+7)]
-----
= g[sqrt(25)]
----
= g(5)
---
= 5*5^2-2*5+3
----
= 125-10+3
---
= 118
================
Cheers,
Stan H.
================